# Coarse-graining particulate two-phase flow

### v7

@@ 4 Conclusions
We have presented a general method to coarse-grain the equations of motion of a mixture of a continuous fluid of arbitrary rheology and non-Brownian particles, interacting via contacts, of arbitrary shape and composition (Supplementary Material), and exemplary applied it to derive the coupled two-fluid mass and momentum balances of a mixture of an incompressible fluid and particles of constant density (§2). The method’s key novelties are as follows: First, the derived micromechanical expressions for the macroscopic fields of interest, such as the effective fluid and solid phase stresses in Eqs. (30) and (37), respectively, are mathematically exact for any ratio between the typical particle size $R$, or the typical particle separation distance, and the macroscopic flow scale $L$, and can be readily extracted from DNS-DEM simulations in a computationally cheap manner (e.g., Appendix C). In contrast, previous exact expressions contain infinite sums of derivatives of ever higher order [12, 13, 14]. Their extraction from DNS-DEM data is not only inconvenient but also computationally costly, since ever finer grids are required to reliably calculate these increasing-order derivatives. Second, the microscopic volume fraction of each particle is its corresponding indicator function, rather than the traditional volume-weighted delta distribution at its center of mass, to ensure that the resulting macroscopic fluid and solid volume fractions add precisely to unity. This leads to an additional contact stress contribution not seen in standard coarse-grained expressions for granular matter (see Eq. (39)), and to particle-rotational contributions to translational macroscopic solid phase balance equations (e.g., see Eq. (38)). As a side note, the fact that these particle-rotational contributions vanish identically for spherical particles suggests that the coarse-graining method may also find application when studying grain shape effects in pure granular flows.
@@
Supplementary information The online version contains supplementary material available at https://xxx.xxx. Furthermore, a MATLAB code calculating the coarse-grained fields from the raw data produced by the IBM-based DNS-DEM simulations of sediment transport is available at https://doi.org/10.5281/zenodo.17499633.
@@ Appendix A Mathematical identities
where $A[\bm{x}]$ is a field well-defined within $\mathbb{V}^{p}$. In particular, the case $A=1$ implies

### v8

@@ 4 Conclusions
We have presented a general method to coarse-grain the equations of motion of a mixture of a continuous fluid of arbitrary rheology and non-Brownian particles, interacting via contacts, of arbitrary shape and composition (Supplementary Material), and exemplary applied it to derive the coupled two-fluid mass and momentum balances of a mixture of an incompressible fluid and particles of constant density (§2). The method’s key novelties are as follows: First, the derived micromechanical expressions for the macroscopic fields of interest, such as the effective fluid and solid phase stresses in Eqs. (30) and (37), respectively, are mathematically exact for any ratio between the typical particle size $R$, or the typical particle separation distance, and the macroscopic flow scale $L$, and can be readily extracted from DNS-DEM simulations in a computationally cheap manner (e.g., Appendix C). In contrast, previous exact expressions contain infinite sums of derivatives of ever higher order [12, 13, 14]. Their extraction from DNS-DEM data is not only inconvenient but also computationally costly, since ever finer grids are required to reliably calculate these increasing-order derivatives. Second, the microscopic volume fraction of each particle is its corresponding indicator function, rather than the traditional volume-weighted delta distribution at its center of mass, to ensure that the resulting macroscopic fluid and solid volume fractions add precisely to unity. This leads to an additional contact stress contribution not seen in standard coarse-grained expressions for granular matter (see Eq. (39)), and to particle-rotational contributions to translational macroscopic solid phase balance equations (e.g., see Eq. (38)). As a side note, the fact that these particle-rotational contributions vanish identically for spherical particles suggests that the coarse-graining method may potentially also find application when studying grain shape effects in pure granular flows.
@@
Supplementary information The online version contains supplementary material available at https://doi.org/10.1007/s10035-025-01609-5.
[…]
Code availability A MATLAB code calculating the coarse-grained fields from the raw data produced by the IBM-based DNS–DEM simulations of sediment transport is available at https://doi.org/10.5281/zenodo.18146169.
@@
[…]
Data availability The data shown in Fig. 3 can be produced using the MATLAB source files uploaded at https://doi.org/10.5281/zenodo.18146169.
@@ Appendix A Mathematical identities
where $A$ is a field well-defined within $\mathbb{V}^{p}$. In particular, the case $A=1$ implies
@@ C.4 Coarse-grained fluid-related quantities
$\begin{split}&\left\langle\sum\nolimits_{p}S^{p}\underline{\bm{r}^{p}\bm{\tilde{f}}^{p}_{\mathrm{IBM}}\delta^{p}_{\mathrm{L}\bm{x}}}_{\mathcal{S}^{p}}\right\rangle_{t}[z_{n}]\\ &\simeq\frac{1}{V_{z}}\sum\nolimits_{p}\sum\nolimits_{l}\bm{r}^{p}_{l}\bm{\tilde{f}}^{p}_{\mathrm{IBM}l}w^{p}_{n}[z_{l}]S^{p}_{l},\end{split}$ (90)